OpenStudy (anonymous):

can anyone tell me how to use algebra to evaluate lim (x+3)^3-27/ x ...which is x goes to 0

OpenStudy (jamesj):

So first of all, simplify that expression you have. Then take the limit.

OpenStudy (asnaseer):

use the formula for the difference between two cubes:\[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (asnaseer):

that will allow you to simplify the expression

OpenStudy (asnaseer):

@hooverst do you understand?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

so what do you get after simplifying?

OpenStudy (asnaseer):

this might help:\[(x+3)^3-27=(x+3)^3-3^3\]

OpenStudy (asnaseer):

now use:\[a^3-b^3=(a-b)(a^2+ab+b^2)\]with:\[a=x+3\]and:\[b=3\]

OpenStudy (anonymous):

\[\Large \lim_{x \rightarrow 0} \frac{(x+3)^3-27}{x}\] \[\Large 27=3^3\] \[\Large \lim_{x \rightarrow 0} \frac{(x+3)^3-3^3}{x}\] Add 3 and -3 to the denominator .... \[\Large \lim_{x \rightarrow 0} \frac{(x+3)^3-3^3}{(x+3)-3}\]

OpenStudy (asnaseer):

@Eyad if you use the formula for the difference between two cubes you will find the x cancels out

OpenStudy (anonymous):

@asnaseer :Ikr,But i liked traditional Law property .

OpenStudy (anonymous):

Since it can be easily formed

OpenStudy (asnaseer):

oh I see what you were trying to do there - yes that would also work

OpenStudy (asnaseer):

@hooverst - have you worked it out yet?

OpenStudy (anonymous):

working on it now

OpenStudy (anonymous):

I'm not sure if it's right but I got 1/-3 but I may have did something wrong.

OpenStudy (asnaseer):

no thats not right - please show your steps so we can help spot where you may have gone wrong.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}(x+3)^{3}-27\div \ x+3-3)\] \[(x+3)^2\left( x+3 \right)-9\div \left( x+3 \right)-3\] \[\left( x+3 \right)^{2}-9\div -3\] \[\left( x ^{2}+9 \right)-9\div -3 =\] \[x ^{2}\div-3=-1/3\]

OpenStudy (asnaseer):

ok, firstly we can use:\[a^3-b^3=(a-b)(a^2+ab+b^2)\]to factor:\[(x+3)^3-27=(x+3)^3-3^3=((x+3)-3)((x+3)^2+3(x+3)+3^2)\]\[\qquad=x((x+3)^2+3(x+3)+9)\]

OpenStudy (asnaseer):

understand so far?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, so we have now shown that:\[(x+3)^3-27=x((x+3)^2+3(x+3)+9)\]therefore, dividing both sides by 'x', we get:\[\frac{(x+3)^3-27}{x}=(x+3)^2+3(x+3)+9\]

OpenStudy (asnaseer):

now you can take the limit of the simplified expression as x tends to zero

OpenStudy (asnaseer):

what answer do you get for this?

OpenStudy (asnaseer):

\[\lim_{x\rightarrow0}\frac{(x+3)^3-27}{x}=\lim_{x\rightarrow0}(x+3)^2+3(x+3)+9\]

OpenStudy (asnaseer):

in the simplified expression, you can just set x=0 and evaluate it

OpenStudy (anonymous):

would be 27

OpenStudy (asnaseer):

that is the correct answer - well done! :)

OpenStudy (anonymous):

wow you are great. I may need your help again.

OpenStudy (asnaseer):

thank you - and you are more than welcome. just post each question in the list to the left and there will be plenty of people here willing to help out.

OpenStudy (anonymous):

ok